# Automated Federated Learning via Informed Pruning

### v1

@@ Abstract
Federated learning (FL) represents a pivotal shift in machine learning (ML) as it enables collaborative training of local ML models coordinated by a central aggregator, all without the need to exchange local data. However, its application on edge devices is hindered by limited computational capabilities and data communication challenges, compounded by the inherent complexity of Deep Learning (DL) models. Model pruning is identified as a key technique for compressing DL models on devices with limited resources. Nonetheless, conventional pruning techniques typically rely on manually crafted heuristics and demand human expertise to achieve a balance between model size, speed, and accuracy, often resulting in sub-optimal solutions.
In this study, we introduce an automated federated learning approach utilizing informed pruning, called AutoFLIP, which dynamically prunes and compresses DL models within both the local clients and the global server. It leverages a federated loss exploration phase to investigate model gradient behavior across diverse datasets and losses, providing insights into parameter significance. Our experiments showcase notable enhancements in scenarios with strong non-IID data, underscoring AutoFLIP’s capacity to tackle computational constraints and achieve superior global convergence.
@@ 1  Introduction
In the past decade, the proliferation of smart devices at the network edge and the surge in data generated and distributed across them have created a decentralized setting. Here, multiple participants store their data locally, which offers an opportunity for collaborative model training, enhancing robustness and generalization. Distributing the computational load across these devices results in faster training times and lower energy consumption compared to centralized approaches.
Nonetheless, distributed machine learning (ML) training encounters significant challenges. One major challenge is ensuring efficient communication and coordination among multiple participants, as each device typically holds only a subset of the data. Effective strategies for aggregating and updating model parameters without excessive communication overhead are crucial. This involves designing algorithms that minimize data exchange while enabling the convergence of high-quality models. Moreover, device heterogeneity, encompassing differences in computational power, storage capacity, and network bandwidth, complicates distributed ML training. Developing algorithms capable of adapting to such environments and utilizing resources efficiently is essential for scaling up distributed learning to large-scale deployments.
@@
Privacy and security concerns are another challenge (Hoffpauir et al., 2023). With sensitive data distributed across various devices, ensuring the privacy of individual data points becomes paramount. For example, medical data stored in hospitals and personal devices is valuable for training diagnostic models but is also subject to strict privacy and security regulations. In this context, Federated Learning  (Zhu et al., 2021) emerges as an effective strategy for training always more complex deep learning (DL) models while preserving the privacy of the data. FL facilitates collaborative model training across multiple devices without exposing local data. A central server, i.e., a global model, coordinates this process by aggregating the updates from locally trained models, which ensures a secure learning environment.
@@
Current FL research primarily focuses on enhancing privacy and adapting ML workflows for specific uses, often with predetermined ML model configurations. For instance, tasks related to computer vision may involve centralized servers selecting well-known neural network architectures like VGG-16 (Simonyan and Zisserman, 2015) (with approximately 138 million parameters) or ResNet-50 (He et al., 2016) (with around 25.6 million parameters). However, the high complexity of these neural networks increases the risk of overfitting, especially with small training data sizes. Additionally, FL systems typically expect clients to have access to high-speed processors and sufficient computational power for local calculations and parameter updates. Yet, many edge devices, such as smartphones, wearables, and sensors, have limited computing and memory capacities, posing a challenge to DL model training systems (Hoffpauir et al., 2023). While DL networks have significantly improved in performance and accuracy, their larger model sizes and increased computational costs present challenges for edge devices that struggle to handle such demands. Moreover, the communication of DL models with millions of parameters presents significant obstacles for FL transmission (Shlezinger et al., 2020; Asad et al., 2023). Therefore, the effective use of FL with edge devices that have limited computational capabilities, while maintaining efficient communication, is an active research question. This underscores the need for personalized and innovative approaches in FL, particularly in optimizing and compressing models to improve inference time, communication cost, energy efficiency, and complexity, all while maintaining a satisfactory accuracy (Rahman et al., 2021).
Numerous methods have been suggested to enhance the effectiveness of neural networks through model compression (Li et al., 2023). Traditionally, this process involves the use of manual rules and specialized expert knowledge to navigate the vast design space, balancing the size, speed and accuracy of the model. However, the vast design space often renders human heuristics a time-consuming strategy for model compression, often leading to suboptimal solutions. Automated Machine Learning (AutoML) systems, like autosklearn (Feurer et al., 2022), TPOT (Olson and Moore, 2016), and H2O AutoML (H2O.ai, 2022), streamline model selection and optimization in centralized settings. Yet, in FL, these systems face challenges due to the distributed nature of data and the impracticality of centralizing it for model selection, given privacy or resource constraints. This backdrop underscores the necessity for novel methods that optimize DL models in FL without centralizing client data, taking into account the available limited resources, the privacy of sensitive information, and addressing class imbalance among the clients in non-IID scenarios.
@@
Our contribution. We introduce a novel automated federated learning approach via informed pruning (AutoFLIP), which uses a loss exploration mechanism (Nikolić et al., 2023) to automatically prune and compress DL models based on shared insights into gradient behaviors across clients. This strategy allows for dynamically reducing the complexity of the DL models in FL environments, hence optimizing performance with limited computational resources at the client level. In our assumed single-server architecture, each client operates on the same initial deep neural network structure that automatically prunes itself at each round, based on the extraction of shared knowledge for an informed model compression. With our experiments over various datasets, tasks, and realistic non-IID scenarios, we provide strong evidence of the effectiveness of the proposed method.
@@
Reproducibility. Our code for reproducing the experiments is available on GitHub.111 https://github.com/ChristianInterno/AutoFLIP
@@ 2.1  Federated Learning
One widely accepted FL standard is known as FedAvg (McMahan et al., 2023). It operates on a distinctive collaborative principle. The global model disseminates a model to different clients for localized training. The impact of each client on the global model update is determined by the size of its dataset, ensuring that clients with larger datasets have a more significant influence on the final global model. This is accomplished by calculating the server parameters as a weighted average of the individual parameters learned by each client. This iterative process, characterized by collaborative learning dynamics, continues until the model reaches an satisfactory termination criterion. Empirical studies have shown the robustness of this approach, even when handling non-convex optimization problems (Das et al., 2022). As a result, it is commonly used as a standard for evaluating newly developed FL protocols. In this study, we will compare the performance of the proposed AutoFLIP method to FedAvg, with and without random pruning, as tested in (Wu et al., 2023). Indeed, among various parameter selection criteria, random pruning demonstrated rapid convergence within their proposed framework, outperforming any other pruning heuristic.
2.2  Model Pruning
Following the assumption that a DL model can contain a sub-network that represents the performance of the entire model after being trained, model pruning is a good strategy to reduce computational requirements of resource-constrained devices. First attempts of using pruning to deploy deep neural networks on resource-limited devices have considered pre-trained convolutional neural networks in a centralized setting (You et al., 2019; Lin et al., 2020). After this, there has also been work focused on dynamic active pruning to increase communication efficiency and reduce the computation cost on the clients during training (Liu et al., 2022; Zhou et al., 2021). Nonetheless, this increases the amount of calculations at the client level. Jiang et al. (2023) introduce PruneFL, a FL method that incorporates adaptive and distributed parameter pruning. Their approach utilizes an unstructured method that does not take advantage of the collective insights of participating clients to develop a cooperative structured pruning strategy. This is in contrast to the objectives of AutoFLIP, which seeks to harness client-specific knowledge to facilitate a structured approach to pruning. Lin et al. (2022) indroduced a novel approach for adaptive per-layer sparsity, however without incorporating any parameter aggregation scheme to reduce the error caused by pruning. This challenge was addressed by Tingting et al. (2023), who were the first applying model pruning directly into a federated learning framework. They moved the pruning process to the global model that works on a computationally more powerful server. The pruned model is distributed to each client, where it undergoes training. Subsequently, each client sends back to the server only the updated parameters, restoring the full structure of the model at the server. Although this study includes various parameter selection criteria from the literature, its pruning method does not incorporate the information gathered during model training. This contrasts with our strategy, AutoFLIP, which leverages such information to enhance the pruning process. The work by Yu et al. (2021) explores enhancing FL with an adaptive pruning method to remove less important parameters, but this adds complexity and overhead. Yu et al. (2023) then proposed Resource-aware Federated Foundation Models, focusing on integrating large transformer-based models into FL, with the limitation of not exploring other architectures. Our method, AutoFLIP, diverges by introducing a pruning strategy that avoids the need for continuous evaluation of parameter significance and is universally applicable across various FL aggregation algorithms and model architectures.
@@ 3  Methodology
AutoFLIP is an automatized FL approach relying on informed pruning. Preliminary step to the FL rounds is an initial exploration phase. Here, a portion or the totality of the clients, which inherit their topology from the global model, are trained for a certain number of exploration epochs. Based on this, for each client, we compute a local guidance matrix, which records how important a certain parameter (weight or bias) is in terms of parameter deviation during the exploration, which reflects loss variability. Afterward, we aggregate the information collected locally in a global guidance matrix, which guides the pruning of the global model.
In each FL round, only a subset of clients influence the global model pruning. The global guidance matrix is recalculated by summing entries from local matrices of selected clients. Using the updated global guidance matrix, the global model pruned parameters are adjusted. Subsequently, the reduced model is distributed to participating clients, who train their pruned models. Server parameters are computed as a weighted average of individual client parameters. The iterative procedure, consisting of (1) pruning of the global and local models, (2) training of the reduced local models, (3) computation of the global model parameters, and (4) local and global model performance evaluation, continues until a prescribed termination criterion is met.
The pruning workflow of AutoFLIP is illustrated in Figure 1. Please note that the initial exploration, computation of weight deviations, and definition of local guidance matrices occur only once. In contrast, the global guidance matrix and subsequent pruning are redefined at each FL round, considering the clients participating in that round.
@@
Figure 1: Illustration of the pruning procedure. The local guidance matrices are computed a priori through an exploration phase. The global guidance matrix is computed aggregating the elements of the local guidance matrices corresponding to the clients participating in each FL round. The global model is pruned and the reduced model is downloaded by the clients.
In the following, we present the notation used to formalize the above procedure and delve deeper into the problem formulation, exploring specific phases of the AutoFLIP workflow.
Notation
We consider a total number of $C$ clients. At each FL round, $K$ clients are chosen and trained on different batches of size $B$ at each epoch, for a total amount of $E$ epochs. The total number of rounds is $R$, which represents our termination criterion. For the exploration phase, we denote with $C_{exp}$ the number of clients selected, which, in this study, we take as the totality $C$ of available clients. The exploration lasts for $E_{exp}$ epochs. For our informed pruning technique, we define a threshold $T_{p}$ for the parameter deviations in the exploration.
Problem Definition and Objective
[…]
Our objective is to automatically determine the best structured pruning strategy to significantly compress client models while aligning the global model parameters $W_{global}$ with the optimal $W_{global}^{*}$ that minimizes the combined loss:
$\min_{h_{i}\in H}\frac{1}{C}\sum_{i=1}^{C}L_{P_{i}}(h_{i},W_{i,t}),$ (1)
where $W_{i,t}$ represents the parameters of client $c_{i}$ at iteration $t$.
Pruning with AutoFLIP
@@
In AutoFLIP, the model initialization phase is augmented by a crucial loss federated exploration phase, allowing clients to explore their loss function landscapes. This exploration phase is completed for embedding gradient variability information into a Pruning Guidance mask $PG_{\text{global}}$. This mask is updated and shared among participating clients in each FL round to guide the evolution of client model structures within an informed pruning session.
@@
To construct the mask $PG_{\text{global}}$, we need an initial exploration phase conducted on $C_{exp}$ clients. In this study, we consider $C_{exp}=C$. We train the clients for $E_{exp}$ epochs, and, for each model parameter, we evaluate its evolution in the search space during the loss exploration phase. This evaluation is conducted by calculating the deviation $D_{1,m}$ for the $m^{th}$ parameter of a client model $c_{i}$ as the squared difference between the initial ($W_{i,m}^{\text{Initial}}$) and final ($W_{i,m}^{\text{Final}}$) parameter values, following $E_{exp}$ epochs of exploration:
@@
Given that we use stochastic gradient descent for the exploration, this deviation acts as a measure of loss variability dependent on the parameter $m$. The greater the variation in the parameter space, the more significant the improvement in loss. The $C_{exp}$ clients compile these deviations into a local matrix $G_{local}$, whose entries are the deviations for the model parameters.
@@
At each FL round, where only $K$ clients are involved, the server aggregates the $G_{local}$ matrices associated to those client to formulate $G_{\text{global}}$ through a normalization process:
$G_{\text{global}}=\frac{1}{K}\sum_{k=1}^{K}\frac{G_{{local}_{k}}-\min(G_{\text{local}})}{\max(G_{\text{local})}-\min(G_{\text{local}})}$ (3)
@@
Each element of $G_{\text{global}}$ thus represents the mean normalized deviation for each parameter, scaled between 0 and 1. A value closer to 0 indicates minimal deviation, suggesting gradient stability during the exploration, hence scarce relevance of the parameter itself. Conversely, values near 1 highlight significant parameter deviations, pointing to more dynamic and potentially insightful areas of the loss landscape. Then, a binarization process is then applied to $G_{\text{global}}$ where elements below $T_{p}$ are set to 0 and those above are set to 1:
$PG_{\text{global},m}=\begin{cases}0&\text{if }G_{\text{global},m}<T_{p}\\ 1&\text{otherwise}\end{cases}$ (4)
@@
Given their smaller influence, parameters corresponding to 0 are flagged for pruning, whereas those marked with 1 are retained, indicating important search directions within the model parameter space. During each FL round, the $K$ participating clients update $PG_{\text{global}}$ by incorporating their $G_{\text{local}}$ deviation values derived from the initial loss exploration phase.
Federated Learning with AutoFLIP
@@
1:Server Initialization: Initial weight matrix $W_{global}^{(0)}$, number of clients for exploration $C_{exp}$, exploration epochs $E_{exp}$, pruning threshold $T_{p}$, FL rounds $R$, training epochs $E$, number of selected clients per round $K$
2:Server selects $C_{exp}$ clients for exploration
3:$G_{{local}_{i}}=(W_{i}^{\text{Initial}}-W_{i}^{\text{Final}})^{2},\forall i\in[1,C_{exp}]$
@@
14:     $W_{global}^{(r)}=\frac{1}{K}\sum_{k=1}^{K}W_{k,\text{pruned}}^{(E)}$
@@
Here our aim is to argument how the parameter pruning mechanism based on loss exploration enters a general FL edge training framework. Algorithm 1 provides an overview of the entire framework of the proposed AutoFLIP algorithm for FL. It is composed by the following steps.

### v2

# Adaptive Hybrid Model Pruning in Federated Learning through Loss Exploration

@@ Abstract
The rapid proliferation of smart devices coupled with the advent of 6G networks has profoundly reshaped the domain of collaborative machine learning. Alongside growing privacy-security concerns in sensitive fields, these developments have positioned federated learning (FL) as a pivotal technology for decentralized model training. Despite its vast potential, specially in the age of complex foundation models, FL encounters challenges such as elevated communication costs, computational constraints, and the complexities of non-IID data distributions. We introduce AutoFLIP, an innovative approach that utilizes a federated loss exploration phase to drive adaptive hybrid pruning, operating in a structured and unstructured way. This innovative mechanism automatically identifies and prunes model substructure by distilling knowledge on model gradients behavior across different non-IID client losses topology, thereby optimizing computational efficiency and enhancing model performance on resource-constrained scenarios. Extensive experiments on various datasets and FL tasks reveal that AutoFLIP not only efficiently accelerates global convergence, but also achieves superior accuracy and robustness compared to traditional methods. On average, AutoFLIP reduces computational overhead by 48.8% and communication costs by 35.5%, while improving global accuracy. By significantly reducing these overheads, AutoFLIP offer the way for efficient FL deployment in real-world applications for a scalable and broad applicability.
@@ 1 Introduction
The proliferation of smart devices at the network edge, coupled with advancements in 6G networks, has created a decentralized setting [31, 47]. Multiple participants store their data locally, which offers an opportunity for collaborative model training, enhancing robustness and generalization. Distributing the computational load across these devices results in faster training times and lower energy consumption compared to centralized approaches [56, 3]. However, collaborative Machine Learning (ML) faces significant challenges [8]. Efficient communication and coordination among participants are crucial, as each device holds only a subset of the data. This requires designing algorithms that minimize data exchange while ensuring high-quality model convergence. Device heterogeneity, including differences in computational power, storage, and bandwidth, further complicates distributed training. Algorithms must adapt to such environments to scale up distributed learning. Privacy and security concerns, along with regulations like the European GDPR [1], the EU AI Act [9] and the U.S. Secure AI [54] Act add another layer of complexity [21]. With sensitive data distributed across various devices, ensuring the privacy of individual data points becomes essential [14]. For example, medical data stored in hospitals and personal devices is valuable for training diagnostic models but is also subject to strict privacy and security regulations [46].
@@
In this context, Federated Learning (FL) [65] emerges as an effective strategy for training always more complex DL models while preserving the privacy of the data. FL facilitates collaborative model training across multiple devices without exposing local data. A central server, i.e., a global model, coordinates this process by aggregating updates from locally trained models, which ensures a secure learning environment. Current FL research focuses on enhancing privacy and adapting ML workflows for specific uses, often with predetermined ML model configurations. Tasks related to computer vision may involve well-known neural network (NN) architectures like VGG-16 [49] (138 million parameters) or ResNet-50 [19] (25.6 million parameters). However, these complex NN risk overfitting, especially with small training data sizes.
@@
In the era of foundation models [3] becoming the norm in machine learning development, FL systems typically expect clients to have high-speed processors and sufficient computational power for local calculations and parameter updates. Yet, many edge devices, such as smartphones, wearable, and sensors, have limited computing and memory capacities, posing a challenge to DL model training systems [21]. Additionally, communicating DL models with millions of parameters presents significant obstacles for FL transmission [48, 2]. Therefore, using FL effectively with edge devices that have limited computational capabilities, while maintaining efficient communication, remains an active research question. FL’s effectiveness is further hindered by the prevalence of non-IID data in real-world scenarios [65, 23, 27]. non-IID data refers to the unique statistical properties of each client’s dataset, reflecting their varied environments. This creates conflicting training goals for local and global models, leading to convergence towards different local optima. As a result, client model updates become biased, impeding global convergence [23]. These challenges underscore the need for personalized and innovative approaches in FL, particularly in optimizing and compressing models to improve inference time, communication cost, energy efficiency, and complexity, all while maintaining satisfactory accuracy.
@@
Our contribution. We introduce a novel automated federated learning approach via adaptive hybrid pruning (AutoFLIP), which uses a novel loss exploration mechanism to automatically prune and compress DL models. In our assumed single-server architecture, each client operates on the same initial deep NN structure that automatically prunes itself at each round, based on the extraction of shared knowledge from the federated loss exploration for an informed model compression. Specifically, by analyzing the variability of gradients during a preliminary local loss exploration phase, which provides insights into gradient behaviors on the loss landscapes across clients, and subsequent information aggregation, the DL models involved in a FL round are hybridly pruned automatically. This strategy allows for dynamically reducing the complexity of the models in FL environments, thereby optimizing performance with limited computational resources at the client level. With our experiments over various datasets, tasks, and realistic non-IID scenarios, we provide strong evidence of the effectiveness and efficiency of AutoFLIP.
@@
Reproducibility. Our code for reproducing the experiments is available on Anonymous GitHub.111 https://github.com/ChristianInterno/AutoFLIP
@@ 2 Background and Related Work
Pruning in Deep Learning. Following the assumption that a DL model can contain a sub-network that represents the performance of the entire model after being trained, model pruning is a good strategy to reduce computational requirements of resource-constrained devices [41, 29, 25]. Most pruning approaches balance accuracy and sparsity during the inference stage by calculating the importance scores of parameters in a well-trained NN and removing those with lower scores. These scores can be derived from weight magnitudes [25, 17], first-order Taylor expansion of the loss function [42, 39], second-order Taylor expansion [29, 18, 40], and other variants [37, 50].
Another recent research direction in NN pruning focuses on improving training efficiency, divided into two categories: pruning at initialization and dynamic sparse training. Pruning at initialization involves pruning the original full-size model before training based on connection sensitivity [30], Hessian-gradient product [55], and synaptic flow [51]. However, since this method does not involve training data, the pruned model may be biased and not specialized for the task. Dynamic sparse training iteratively adjusts the pruned model structure during training while maintaining the desired sparsity [7, 10]. This approach requires memory-intensive operations due to the large search space, making it impractical for resource-constrained devices.
Initial attempts to use pruning for deploying deep neural networks on resource-limited devices have utilized pre-trained CNNs in a centralized setting [61, 33]. However, this approach can lead to reduced data privacy, higher costs, poor adaptation to local conditions, suboptimal performance on diverse data, and latency in real-time applications.
Hybrid Pruning in Deep Learning. Hybrid pruning techniques combine structured and unstructured pruning strategies to optimize both performance and efficiency in deep neural networks [36, 44, 13, 15]. Structured pruning [20] removes entire units like neurons, filters, or layers, leading to more hardware-efficient designs that are easier to implement on resource-constrained devices. On the other hand, unstructured pruning [32]focuses on removing individual weights, which can achieve higher sparsity levels and further compress the model, though it may require more complex hardware support [40].
Pruning in Federated Learning. The widely accepted FL standard is known as FedAvg [38]. It distributes a global model to clients for local training and aggregates it by averaging their parameters. Empirical studies have shown the robustness of this approach, even when handling non-convex optimization problems [5]. As a result, it is commonly used as a standard for evaluating newly developed FL protocols. In this study, we will compare the performance of the proposed AutoFLIP method to FedAvg, with different State-of-the-Art (SotA) FL pruning approachs, as tested in [58]. In fact, since data remains locally stored and cannot be shared, traditional centralized pruning approaches that rely on access to training data are not feasible in FL.
In the context of FL, there has been work focused on dynamic active pruning to increase communication efficiency during training. Liu et al. [35], Zhou et al. [64] introduced a method where pruning decisions are made dynamically based on the model’s real-time performance evaluation, which significantly reduces the data exchanged during training but adds computational complexity to client devices. Jiang et al. [26] introduced PruneFL, a FL method that incorporates adaptive and distributed parameter pruning. Their approach utilizes an unstructured method that does not take advantage of the collective insights of participating clients to develop a cooperative structured pruning strategy. This is in contrast to the objectives of AutoFLIP, which seeks to harness client-specific knowledge to facilitate a structured approach to pruning. Lin et al. [34] introduced a novel approach for adaptive per-layer sparsity, however without incorporating any parameter aggregation scheme to reduce the error caused by pruning. This challenge was addressed by Tingting et al. [52] by moving the pruning process to the global model that works on a computationally more powerful server. The pruned model is distributed to each client, where it undergoes training. Subsequently, each client sends back to the server only the updated parameters, restoring the full structure of the model at the server. Although this study includes various parameter selection criteria from the literature, its pruning method does not incorporate the information gathered during model training. This contrasts with our strategy, AutoFLIP, which leverages such information to enhance the pruning process. Yu et al. [62] proposed Resource-aware Federated Foundation Models, focusing on integrating large transformer-based models into FL, with the limitation of not exploring other architectures. Our method, AutoFLIP, diverges by introducing a pruning strategy that avoids the need for continuous evaluation of parameter significance and is universally applicable across various FL aggregation algorithms and model architectures.
3 Preliminaries
Notation: We consider a total number of $C$ clients. At each FL round, $K$ clients are chosen and trained on different batches of size $B$ for $E$ epochs. The total number of rounds is $R$, which represents our termination criterion. For the exploration phase,$C_{\text{exp}}$ is the number of clients selected, which, in this study, we take as the totality $C$ of available clients. The exploration lasts for $E_{\text{exp}}$ epochs.
@@ 3.1 Federated Learning
In the conventional FL setting, each client $i$ ($1\leq i\leq K$) possesses its own data distribution $p_{i}(x,y)$, where $x\in\mathbb{R}^{d}$ represents the $d$-dimensional input vector and $y\in\{1,\ldots,M\}$ is the corresponding label from $M$ classes. Each client has a dataset $D_{i}$ with $N_{i}$ data points: $D_{i}=\{(x_{i}^{(1)},y_{i}^{(1)}),\ldots,(x_{i}^{(N_{i})},y_{i}^{(N_{i})})\}$. It is assumed that in a non-IID scenario the data distribution $p_{i}(x,y)$ varies across clients. These data distributions $p_{i}(x,y)$ are sampled from a family $\mathcal{E}$ of distributions. The objective is for the clients to collaboratively train a global model with parameters $W_{\text{global}}$, which will perform predictions on new data. The global loss function for a data point $(x,y)$ is denoted by $\mathcal{L}(W_{\text{global}},x,y)$, where the global objective function to be minimized is defined as: $\mathcal{L}(W_{\text{global}}):=\frac{1}{C}\sum_{i=1}^{C}\mathbb{E}_{(x_{i},y_{i})\sim p_{i}}[\mathcal{L}(W_{\text{global}},x_{i},y_{i})],$ with $\mathbb{E}_{(x_{i},y_{i})\sim p_{i}}$ representing the expected loss over the data distribution $p_{i}$ for each client $i$ with parameters $W$.
The optimization process involves several key steps:
1. Client Selection: A subset of $K$ clients is selected from the total $C$ clients. 2. Local Training: Each selected client $i$ performs local training for $E$ epochs using its local dataset $D_{i}$. The local training aims to minimize the local objective function $\mathcal{L}(W_{i})$ using stochastic gradient descent (SGD): let $W_{i}^{r}$ be the local model parameters of client $i$ at round $r$, the update rule is given by: $W_{i}^{r+1}=W_{i}^{r}-\eta\nabla\mathcal{L}(W_{i}^{r})$, where $\eta$ is the learning rate. 3. Parameter Aggregation: After local training, each client sends its updated parameters $W_{i}^{r+1}$ to the central server. The server aggregates these parameters to form the new global model $W_{global}^{r+1}$ using a weighted average: $W_{global}^{r+1}=\frac{1}{K}\sum_{i=1}^{K}W_{i}^{r+1}$. This iterative process is repeated for $R$ rounds the termination criterion is met.
3.2 Problem Definition and Objective
i) Mitigates noise and biases in the client trajectories: By selectively pruning model parameters based on their contribution to loss topology variability, we aim to the trajectories of the different clients to help converging the global model more effectively despite the heterogeneous data distributions across clients. ii) Enhance Computation and Communication Efficiency: The hybrid pruning serves as a mechanism to decrease the number of model parameters that need to be communicated between clients and the central server. By eliminating less critical parameters and substructure, AutoFLIP reduces communication overhead and accelerates the computational process. This not only expedites the overall FL workflow but also makes it more feasible to deploy in resource-constrained scenarios.
@@ 4 Methodology
AutoFLIP is an automatized FL approach that utilizes informed pruning through a federated client loss exploration process. Inspired by the idea of utilizing agents with similar tasks as scouts which explore the conformation of different loss function landscapes from Internò et al. [23], Nikolić et al. [43], AutoFLIP introduces a preliminary step to the FL process, which we term federated loss exploration phase. Here, a $C_{\text{exp}}$ portion of clients (or the totality $C$), which inherit their model structure from the global model, explore for a number of $E_{\text{exp}}$ exploration epochs their loss landscape using its local dataset $D_{i}$. Based on this, for each client $c_{\text{exp}_{i}}$, we compute a local guidance matrix $G_{\text{local}_{i}}$, which records how important a certain parameter $W_{i}$ (weight or bias) is in terms of loss variability 1. Afterward, we aggregate the information collected locally in a global pruning guidance matrix $PG_{\text{global}}$ on the server 2, which will generate an informed pruning mask to guide the pruning of the client models 3. The pruning workflow of AutoFLIP is illustrated in Figure 1. Please note that the initial federated loss exploration, computation of parameter deviations, and definition of local guidance matrices occur only once at the beginning of the FL optimization process as a preliminary procedure. In contrast, the global guidance matrix and subsequent pruning strategy are automatically redefined in each FL round, considering the clients participating in that round.
To summarize, the iterative procedure consists of (1) pruning local models using the updated pruning guidance matrix, (2) training the pruned local models, (3) aggregating the model parameters and (4) evaluating performance and updating the pruning guidance matrix in each FL round.
@@
Figure 1: Illustration of the AutoFLIP pruning procedure. The local guidance matrices are computed a priori through the federated exploration phase. The global guidance matrix is computed by the server by aggregating the elements of the local guidance matrices corresponding to the clients participating in each FL round. The pruning mask is downloaded by the participant clients. All steps preliminary to the FL procedure are denoted in gray, while the steps intrinsic to the FL procedure with pruning are denoted in red.
4.1 Federated Loss Exploration
@@
In AutoFLIP, the model initialization phase is augmented by a crucial federated loss exploration phase, allowing clients to explore their loss function landscapes. We envision each client as an explorer that delves into different regions of their loss landscape. Through this exploration, they can identify crucial dimensions and those that can be disregarded based on their experience by quantifying gradient variability during the exploration. In other words, how much the loss is steep in that direction. Subsequently, they transmit this knowledge to the server, which updates a pruning guidance mask $PG_{\text{global}}$. This knowledge contained in the mask is then distilled among participating clients in each FL round to guide the evolution of client model structures within an informed pruning session.
@@
To construct the mask $PG_{\text{global}}$, we begin with an initial exploration phase conducted on $C_{\text{exp}}$ clients. In this study, we consider $C_{\text{exp}}=C$. In our study, we let explore the clients for $E_{\text{exp}}=150$ epochs. Appendix A provides ablation studies for $C_{\text{exp}}$ and $E_{\text{exp}}$. For each model parameter we evaluate its evolution in the search space during the loss exploration. This evaluation is conducted by calculating the deviation $D_{i,m}$ for the $m^{th}$ parameter of a client model $i$ as the squared difference between the initial ($W_{i,m}^{\text{Initial}}$) and final ($W_{i,m}^{\text{Final}}$) parameter values after $E_{\text{exp}}$ epochs of exploration:
@@
Using stochastic gradient descent for exploration, the deviation $D_{i,m}$ in Eq. (1) serves as a measure of gradient variability on the loss landscape for parameter $m$ during the preliminary exploration phase before the actual FL procedure. The greater the variation in the parameter space, the faster the improvements in loss: the update rule for a parameter in stochastic gradient descent is $W_{{i,m}}^{(e_{\text{exp}}+1)}=W_{{i,m}}^{(e_{\text{exp}})}-\eta\nabla\mathcal{L}_{i}\left(W_{{i,m}}^{(e_{\text{exp}})};D_{i}\right))$, where $W_{{i,m}}^{(e_{\text{exp}})}$ and $W_{{i,m}}^{(e_{\text{exp}}+1)}$ are the values of the parameter $m$ at the exploration epochs $e_{\text{exp}}$ and $e_{\text{exp}}+1$, $\eta$ is the learning rate, and $\nabla\mathcal{L}_{i}\left(W_{{i,m}}^{(e_{\text{exp}})}\right)$ is the gradient of the loss function of client $i$ with respect to the parameter $m$ at epoch $e_{\text{exp}}$ using its local dataset $D_{i}$. Given the gradient update rule, the deviation in $W_{i,m}$ from the initial to the final exploration epoch $E_{\text{exp}}$ can be approximated to $W_{i,m}^{\text{Final}}-W_{i,m}^{\text{Initial}}\approx-\eta\sum_{t=1}^{E_{\text{exp}}}\nabla\mathcal{L}_{i}\left(W_{i,m}^{(t)};D_{i}\right)$. To ensure non-negativity and highlight larger deviations more severely, we take the square of this value. This squared deviation measure $D_{i,m}$ approximates the square of the sum of gradients affecting the parameter evolution, indicating the significance of parameter updates on loss variability during the exploration phase. By squaring the sum of the gradients, we ensure that the deviation measure is always non-negative and that larger deviations are highlighted more severely than smaller ones.
@@
The $C_{\text{exp}}$ clients compile these deviations into a local matrix $G_{\text{local}}$, whose entries are the deviations for the model parameters. At each FL round, where only $K$ clients are involved, the server aggregates the $G_{\text{local}}$ matrices associated to those client to formulate $G_{\text{global}}$ through a normalization process:
$G_{\text{global}}=\frac{1}{K}\sum_{k=1}^{K}\frac{G_{\text{local}_{k}}-\min(G_{\text{local}})}{\max(G_{\text{local})}-\min(G_{\text{local}})}$ (2)
@@
Here, the minimum and maximum values are taken over all $G_{\text{local}_{k}}$ matrices for $k=1,\ldots,K$. Each element of $G_{\text{global}}$ thus represents the mean normalized deviation for each parameter, scaled between 0 and 1. This process ensures that no single client’s $G_{\text{local}}$ disproportionately influences $G_{\text{global}}$ due the possible presence of outliers in terms of deviations $D_{i,m}$. A value closer to 0 indicates minimal deviation, suggesting gradient stability during the exploration, hence scarce relevance of the parameter itself. Conversely, values near 1 highlight significant parameter deviations, pointing to more dynamic and potentially insightful areas of the loss landscape. Then, a binarization process is applied to $G_{\text{global}}$ where elements below $T_{\text{p}}$ are set to 0 and those above are set to 1:
$PG_{\text{global},m}=\begin{cases}0&\text{if }G_{\text{global},m}<T_{\text{p}}\\ 1&\text{otherwise}\end{cases}$ (3)
@@
The threshold $T_{\text{p}}$ directly determines the compression ratio of the model by setting the proportion of parameters to be pruned. $T_{\text{p}}$ can be seen has a hard constraint on model size, which is sometimes needed due to devices’ resource constraints. Given their smaller influence, parameters corresponding to 0 are marked for pruning, whereas those marked with 1 are retained, indicating important search directions within the model parameter space. During each FL round, the $K$ participating clients update $PG_{\text{global}}$ by incorporating their $G_{\text{local}}$ deviation values derived from the initial loss exploration phase.
To select an appropriate $T_{\text{p}}$, consider the desired compression ratio for the model. This ratio reflects the extent to which the model needs to be compressed while maintaining acceptable performance. By carefully selecting $T_{\text{p}}$ based on the desired compression ratio and empirical validation, we can achieve a well-balanced model that is both efficient and accurate, tailored to the specific needs of the FL task. In appendix B we conduct ablation studies on $T_{\text{p}}$.
The Proposed AutoFLIP Framework. Here our aim is to argument how the parameter pruning mechanism based on loss exploration enters a general FL edge training framework. Algorithm 1 provides an overview of the entire framework of the proposed AutoFLIP algorithm for FL. It is composed by the following steps.
@@
1:Server Initialization: Initial matrix $W_{\text{global}}^{(0)}$, number of clients for exploration $C_{\text{exp}}$, exploration epochs $E_{\text{exp}}$, pruning threshold $T_{\text{p}}$, FL rounds $R$, training epochs $E$, number of selected clients per round $K$
2:Server selects $C_{\text{exp}}$ clients for exploration
3:$G_{\text{local}_{i}}=(W_{i}^{\text{Initial}}-W_{i}^{\text{Final}})^{2},\forall i\in[1,C_{\text{exp}}]$
@@
14:     $W_{\text{global}}^{(r)}=\frac{1}{K}\sum_{k=1}^{K}W_{k,\text{pruned}}^{(E)}$triangleright This can be replaced with other FL aggregation algorithms